# Diagnosing and Mitigating Modality Interference in Multimodal Large Language Models

### v3

@@ C.4 Evaluating Chain-of-Thought Prompting for Modality Interference Mitigation
To further investigate the potential of prompt-based methods in mitigating modality interference, we conduct additional experiments using Chain-of-Thought Wei et al. (2022) style prompting. This approach aims to encourage structured reasoning by guiding the model through an explicit reasoning process before producing its final answer.
@@ C.5 Evaluation on Free-Form VQA
To further assess the generalizability of our method beyond multiple-choice VQA tasks, we evaluate on TextVQA (Singh et al., 2019), a free-form generative visual question answering dataset that requires reasoning over both textual and visual content in natural images. We follow the benchmark adopted by LLaVA Liu et al. (2023b), which evaluates a model’s ability to both recognize textual characters within images and effectively handle noisy outputs generated by OCR systems.
@@ C.6 Improving Generalizability through Adversarial Training
  • Document OCR noise: Real-world noisy OCR snippets are sampled from the FUNSD dataset (Jaume et al., 2019) and inserted as irrelevant textual distractors into visual classification tasks (Mini-ImageNet, Caltech-101).
  • Unrelated screenshots: Unrelated UI screenshots are drawn from the RICO dataset (Deka et al., 2017) and added as visual distractors to language-dominant VQA tasks (OpenBookQA, MMLU).
Each experiment is repeated multiple times, and we report average accuracy across runs (standard deviation << 0.1). Results are presented in Table 11. Across all model scales and task types, adversarial training consistently improves robustness to both types of perturbations. These findings indicate that the additional training overhead introduced by adversarial perturbation is well-justified by the improved generalization to unseen distribution shifts—a desirable property for reliable deployment in real-world settings.
@@ Appendix F Limitations
Appendix G Broader Impacts
This work investigates the limitations of current multimodal large language models in reasoning across modalities and proposes methods to mitigate modality interference—a concrete failure case of cross-modality competency. By improving the robustness and alignment behavior of MLLMs, our approach may benefit a variety of downstream applications that rely on accurate visual-linguistic understanding, including education, accessibility tools (e.g., visual question answering for blind users), and scientific multimodal reasoning tasks.
Our findings also highlight the hidden risks of over-relying on irrelevant modality signals, which can degrade performance or lead to misleading predictions. Making such failure modes measurable and diagnosable can support safer deployment and more transparent evaluation of MLLMs in practice.
On the other hand, techniques such as adversarial perturbation may be dual-use. While our implementation uses perturbations to improve model alignment, similar strategies could be misused to manipulate model behavior. To reduce such risks, we restrict all experiments to open-source academic models and do not include harmful or sensitive content in training or evaluation. We encourage future work to further assess modality interference in safety-critical contexts and to investigate alignment-aware perturbation techniques with explicit safety constraints.
Appendix H The Use of Large Language Models

### v4

@@ C.4 Evaluating Chain-of-Thought Prompting for Modality Interference Mitigation
To further investigate the potential of prompt-based methods in mitigating modality interference, we conduct additional experiments using Chain-of-Thought (Wei et al., 2022) style prompting. This approach aims to encourage structured reasoning by guiding the model through an explicit reasoning process before producing its final answer.
@@ C.5 Evaluation on Free-Form VQA
To further assess the generalizability of our method beyond multiple-choice VQA tasks, we evaluate on TextVQA (Singh et al., 2019), a free-form generative visual question answering dataset that requires reasoning over both textual and visual content in natural images. We follow the benchmark adopted by LLaVA (Liu et al., 2023b), which evaluates a model’s ability to both recognize textual characters within images and effectively handle noisy outputs generated by OCR systems.
@@ C.6 Improving Generalizability through Adversarial Training
  • Document OCR noise: Real-world noisy OCR snippets are sampled from the FUNSD dataset (Jaume et al., 2019) and inserted as irrelevant textual distractors into image-heavy visual classification tasks (Mini-ImageNet, Caltech-101).
  • Unrelated screenshots: Unrelated UI screenshots are drawn from the RICO dataset (Deka et al., 2017) and added as visual distractors to text-heavy reasoning tasks (OpenBookQA, MMLU).
Beyond these seen datasets, we further evaluate generalization on two completely unseen datasets at test time: Food-101 and ARC Challenge. Food-101 serves as an unseen image-heavy classification benchmark, where models are exposed to diverse textual perturbations, including unrelated descriptions, misleading class cues, and OCR-style noise. ARC Challenge represents an unseen text-heavy reasoning task, where various visual distractors—such as random pixels, unrelated real images, full-black/white images, and UI screenshots—are introduced to probe robustness against spurious visual signals.
Each experiment is repeated multiple times, and we report average accuracy across runs (standard deviation <0.1<0.1). Results on seen datasets are presented in Table Table 11, while results on unseen datasets are summarized in  Table 12. Across all model families and task types, adversarial training consistently improves robustness to both unseen perturbation types and unseen datasets. In contrast, heuristic data augmentation alone often fails to generalize beyond the perturbation patterns observed during training and can even degrade performance under novel noise.
Overall, these results demonstrate that adversarial training substantially enhances out-of-domain generalization with only modest additional training overhead. This improved robustness to distribution shifts is crucial for reliable deployment of multimodal models in real-world, noisy environments.
@@
Table 12: Accuracy (%) on Food-101 (image-heavy) and ARC Challenge (text-heavy) under diverse test-time perturbations. For Food-101, we add different types of textual distractors (unrelated, misleading, and OCR-style noise). For ARC Challenge, we use various visual distractors (random pixels, unrelated real images, full-black/white images, and UI screenshots).
Setting Model Food-101 ARC Challenge
Origin Unrelated Text Mislead Text OCR Text RandPixels Real Image Full Black Full White Screenshot
vanilla qwen2.5-vl-7b 97.1 96.8 93.3 96.7 85.6 83.7 85.9 85.9 79.6
$D^{\text{AUG}}$ qwen2.5-vl-7b 97.1 96.6 94.1 96.5 85.0 82.9 84.9 84.9 80.2
Ours qwen2.5-vl-7b 97.5 97.5 97.6 97.4 84.1 82.7 84.2 84.0 82.6
vanilla llava-1.5-7b 90.2 86.3 31.0 83.6 53.4 51.6 54.5 54.0 51.8
$D^{\text{AUG}}$ llava-1.5-7b 91.3 86.5 40.2 83.5 55.7 52.0 56.0 56.0 53.8
Ours llava-1.5-7b 92.1 89.7 69.0 89.5 62.8 62.0 62.6 62.8 62.6
vanilla instructblip-vicuna-7b 67.9 20.1 18.7 48.0 33.7 35.8 33.7 31.2 36.5
$D^{\text{AUG}}$ instructblip-vicuna-7b 55.5 41.4 38.4 56.5 27.3 28.2 27.6 27.2 30.7
Ours instructblip-vicuna-7b 95.6 89.8 93.7 88.1 57.6 56.4 58.8 59.6 57.0
C.7 Ablation without Modality Mask
To further examine the role of modality-aware masking, we implement a No Mask variant, where adversarial perturbations are applied uniformly to all modalities across all datasets, including VQA tasks.
Table 13 reports the results on representative image-heavy and text-heavy benchmarks. We observe three key findings. First, training without a modality mask still partially mitigates modality interference. This is because the combination of consistency regularization and heuristic perturbations already encourages the model to learn certain modality-invariant features, and adversarial training itself improves generic robustness (e.g., on OpenBookQA with screenshot distractors, accuracy improves from 50.2% to 62.7%).
Second, the modality mask is critical for out-of-distribution generalization. Without masking, performance collapses on unseen perturbations, particularly under strong visual distractions. For example, on OpenBookQA with screenshot perturbations, the No Mask variant underperforms the masked version by a large margin (62.7% vs. 81.8%).
Third, modality-aware masking is essential to preserve multimodal semantics during VQA training. When adversarial noise is injected into all token embeddings indiscriminately, the model fails to maintain coherent cross-modal representations, resulting in a notable degradation in overall VQA performance (61.5% vs. 66.8%).
Together, these results demonstrate that while adversarial training alone improves robustness, modality-aware masking plays a crucial role in preventing semantic destruction and enabling reliable generalization under unseen distribution shifts.
Table 13: Ablation study without modality-aware masking. The No Mask variant applies adversarial perturbations to all modalities across all datasets. Results are reported on image-heavy (Caltech-101) and text-heavy (OpenBookQA) benchmarks, together with overall VQA accuracy.
Model Caltech-101 OpenBookQA VQA
Orig Mislead Text OCR Text RandPixels Real Image Screenshot Overall
LLaVA-1.5-7B (Vanilla) 97.0 57.4 92.8 62.4 56.4 50.2 63.8
LLaVA-1.5-7B (Ours, No Mask) 98.1 96.3 93.5 79.5 78.1 62.7 61.5
LLaVA-1.5-7B (Ours) 99.3 98.9 99.0 81.8 81.0 81.8 66.8
C.8 Evaluation on Reasoning-Enhanced (“Thinking”) Models
Reasoning-enhanced “thinking” models may potentially mitigate modality interference by explicitly reasoning over multimodal inputs. To partially examine this hypothesis, we first evaluate chain-of-thought (CoT) prompting in the main paper (Table 1). While CoT prompting reliably improves reasoning accuracy on clean inputs, it does not reduce modality interference under modality-mismatched perturbations, and the same failure modes persist.
To further study this question, we additionally evaluate R1-OneVision-7B (Yang et al., 2025), a multimodal large language model explicitly designed with enhanced reasoning capabilities with a backbone Qwen2.5-VL-7B model. We compare its performance against a strong non-reasoning baseline, Qwen2.5-VL-7B, under representative image-heavy and text-heavy OOD settings.
Table 14 reports the results. Despite its stronger reasoning ability, R1-OneVision-7B still exhibits substantial performance degradation under misleading textual perturbations and visual distractors. In contrast, Qwen2.5-VL-7B shows comparable or stronger robustness under the same conditions.
These results suggest that improved reasoning alone is insufficient to resolve modality interference. Without explicit mechanisms to control modality reliance, such as modality-aware perturbation and regularization, even reasoning-enhanced models remain vulnerable to task-irrelevant signals.
We investigated the generated reasoning traces: In some cases the model shows better awareness of irrelevant modalities, but in many cases the model still treats unrelated modality as meaningful evidence and produces long hallucinated interpretations that directly influence the final answer. This confirms that our methods remain necessary for interference mitigation.
Table 14: Evaluation of reasoning-enhanced “thinking” models under modality-mismatched perturbations. Results are reported on image-heavy (Caltech-101) and text-heavy (OpenBookQA) tasks.
Model Caltech-101 (Orig) Caltech-101 (Mislead Text) OpenBookQA (RandPixels) OpenBookQA (Real Image)
R1-OneVision-7B 96.5 84.7 81.2 77.6
Qwen2.5-VL-7B 99.1 97.2 85.9 77.5
@@ Appendix F Limitations
Appendix G The Use of Large Language Models